# Strong order-one convergence of the Euler method for random ordinary differential equations driven by semi-martingale noises

### v7

@@ 8.3.3. Fractional Brownian motion noise
$\begin{cases}\displaystyle\frac{\mathrm{d}X_{t}}{\mathrm{d}t}=-X_{t}+B^{H}_{t},\qquad 0\leq t\leq T,\\ \left.X_{t}\right|_{t=0}=X_{0},\end{cases}$ (8.11)
and where the noise $\{B^{H}_{t}\}_{t}$ is assumed to be a fractional Brownian motion (fBm) with Hurst parameter $10<H<$ 1. It turns out that, for $20<H<1/$ 2, the order of convergence is $H+1/2,$ while for $1/2\leq H<1,$ the order of convergence is 1. In any case, this is higher than the order $H$ estimated in previous works. The same seems to hold for a nonlinear dependency on the fBm, but the proof is more involved, depending on a fractional Itô formula (see [7, Theorem 4.2.6], [4, Theorem 4.1]and [29, Theorem 2.7.4]), based on the Wick Itô Skorohod (WIS) integral (see [7, Chapter 4]). A corresponding WIS isometry is also needed (see e.g. [7, Theorem 4.5.6]), involving Malliavin calculus and fractional derivatives. For these reasons, we leave the nonlinear case to a subsequent work and focus on this simple linear example, which suffices to illustrate the peculiarity of the dependence on $H$ of the order of convergence. For this linear equation, the proof of convergence is done rigorously below, with the framework developed in the main article.
We need to estimate the last term in Equation 6.2 of Proposition 6.1, involving the steps of the term $f(t,x,y)=-x+y$, which in this case reduce to
$f(s,X_{\tau^{N}(s)}^{N},Y_{s})-f(\tau^{N}(s),X_{\tau^{N}(s)}^{N},Y_{\tau^{N}(s)})=B^{H}_{s}-B^{H}_{\tau^{N}(s)},$ (8.12)
for $0\leq s\leq T$. There are several ways to represent an fBm (see e.g. [7, 29]). We use the formula (see [26, eq. (2.1)] or [7, eq. (1.1)])
$B^{H}_{t}=\frac{1}{\Gamma(H+1/2)}\left(\int_{-\infty}^{0}\left((t-s)^{H-1/2}-(-s)^{H-1/2}\right)\;\mathrm{d}W_{s}\right.\\ \left.+\int_{0}^{t}(t-s)^{H-1/2}\;\mathrm{d}W_{s}\right),$ (8.13)
where $\Gamma(\cdot)$ is the well-known Gamma function. For the step, Equation 8.13 means that
$B^{H}_{s}-B^{H}_{\tau^{N}(s)}=\frac{1}{\Gamma(H+1/2)}\left(\int_{-\infty}^{\tau^{N}(s)}\left((s-\xi)^{H-1/2}-(\tau^{N}(s)-\xi)^{H-1/2}\right)\;\mathrm{d}W_{\xi}\right.\\ \left.+\int_{\tau^{N}(s)}^{s}(s-\xi)^{H-1/2}\;\mathrm{d}W_{\xi}\right).$ (8.14)
Then, using the stochastic Fubini Theorem to exchange the order of integration (see, again, [32, Section IV.6]),
@@
$\frac{1}{\Gamma(H+1/2)}\Delta t_{N}\mathbb{E}\left[\left|\int_{-\infty}^{0}\int_{0}^{t_{j}}\left((s-\xi)^{H-1/2}-(\tau^{N}(s)-\xi)^{H-1/2}\right)\;\mathrm{d}s\;\mathrm{d}W_{\xi}\right|\right]\\ \leq c_{H}^{(1)}\Delta t_{N},$ (8.16)
@@
$\frac{1}{\Gamma(H+1/2)}\mathbb{E}\left[\left|\int_{0}^{t_{j}}\int_{\tau^{N}(\xi)+\Delta t_{N}}^{t_{j}}\left((s-\xi)^{H-1/2}-(\tau^{N}(s)-\xi)^{H-1/2}\right)\;\mathrm{d}s\;\mathrm{d}W_{\xi}\right|\right]\\ \leq c_{H}^{(2)}\Delta t_{N},$ (8.17)
@@
$\frac{1}{\Gamma(H+1/2)}\mathbb{E}\left[\left|\int_{0}^{t_{j}}\int_{\xi}^{\tau^{N}(\xi)+\Delta t_{N}}(s-\xi)^{H-1/2}\;\mathrm{d}s\;\mathrm{d}W_{\xi}\right|\right]\leq c_{H}^{(3)}\Delta t_{N}^{H+1/2},$ (8.18)
@@
Putting the three estimates Equation 8.16, Equation 8.17, Equation 8.18 in Equation 8.15 we find that
$\mathbb{E}\left[\left|\int_{0}^{t_{j}}\left(f(s,X_{\tau^{N}(s)}^{N},Y_{s})-f(\tau^{N}(s),X_{\tau^{N}(s)}^{N},Y_{\tau^{N}(s)})\right)\;\mathrm{d}s\right|\right]\\ \leq c_{H}^{(4)}\Delta t_{N}+c_{H}^{(3)}\Delta t_{N}^{H+1/2},$ (8.19)
where $c_{H}^{(4)}=c_{H}^{(1)}+c_{H}^{(2)}$. Using this estimate in Proposition 6.1 shows that the Euler method is of order $H+1/2$, when $20<H<1/$ 2, and is of order 1, when $11/2\leq H<$ 1, having in mind that $H=1/2$ corresponds to the classical Wiener process.
@@
Theorem 8.1.
Consider the equation Equation 8.11 where $\{B^{H}_{t}\}_{t}$ is a fractional Brownian motion (fBm) with Hurst parameter $10<H<$ 1. Suppose the initial condition $X_{0}$ satisfies
$\mathbb{E}[\|X_{0}\|^{2}]<\infty.$ (8.20)
Then, the Euler scheme for this initial value problem is of strong order $H+1/2$, for $20<H<1/$ 2, and is of order $1$, for $11/2\leq H<$ 1. More precisely,
$\max_{j=0,\ldots,N}\mathbb{E}\left[\left|X_{t_{j}}-X_{t_{j}}^{N}\right|\right]\leq c_{1}\Delta t_{N}+c_{2}\Delta t_{N}^{H+1/2},\qquad\forall N\in\mathbb{N},$ (8.21)
@@
The fractional Browninan motion noise term is generated with the $\mathcal{O}(N\log N)$ fast Fourier transform (FFT) method of Davies and Harte, as presented in [13] (see also [18, Section 14.4]). Table 4 shows the obtained convergence estimates, for a number of Hurst parameters, which is illustrated in Figure 12, matching the theoretical estimate of $p=\min\{H+1/2,1\}.$
@@
8.3.4. Population dynamics with harvest
Here, we consider a population dynamics problem modeled by a logistic equation with random coefficients, loosely inspired by [18, Section 15.2], with an extra term representing harvest:
$\frac{\mathrm{d}X_{t}}{\mathrm{d}t}=\Lambda_{t}X_{t}(1-\frac{X_{t}}{r})-\alpha H_{t}\frac{X_{t}}{r+X_{t}},$ (8.23)
@@ 8.3.4. Population dynamics with harvest
$0<\gamma-\varepsilon\leq\Lambda_{t}\leq\gamma+\varepsilon<2\gamma,\qquad\forall t\geq 0$
@@
The equation Equation 8.23 becomes
@@
Table 5. Mesh points (N), time steps (dt), strong error (error), and standard error (std err) of the Euler method for population dynamics for each mesh resolution $N$, with initial condition $X_{0}\sim\mathrm{Beta}(7.0,5.0)$ and gBm and step process noises, on the time interval $I=[0.0,1.0]$, based on $M=100$ sample paths for each fixed time step, with the target solution calculated with $262144$ points. The order of strong convergence is estimated to be $p=1.009$, with the 95% confidence interval $[0.9603,1.0571]$.
Figure 13. Order of convergence of the strong error of the Euler method for the population dynamics equation Equation 8.23, based on Table 5.
Figure 14. An approximation sequence of a sample path solution of the population dynamics equation Equation 8.23.
8.3.5. A toggle-switch model for gene expression
@@
$\begin{cases}\frac{\displaystyle\mathrm{d}X_{t}}{\displaystyle\mathrm{d}t}=\left(A_{t}+\frac{\displaystyle X_{t}^{4}}{\displaystyle a^{4}+X_{t}^{4}}\right)\left(\frac{\displaystyle b^{4}}{\displaystyle b^{4}+Y_{t}^{4}}\right)-\mu X_{t},\\ \frac{\displaystyle\mathrm{d}Y_{t}}{\displaystyle\mathrm{d}t}=\left(B_{t}+\frac{\displaystyle Y_{t}^{4}}{\displaystyle c^{4}+Y_{t}^{4}}\right)\left(\frac{\displaystyle d^{4}}{\displaystyle d^{4}+X_{t}^{4}}\right)-\nu Y_{t},\end{cases}$ (8.24)
@@
Table 6. Mesh points (N), time steps (dt), strong error (error), and standard error (std err) of the Euler method for a toggle-switch model of gene regulation for each mesh resolution $N$, with initial condition $X_{0}=4.0,Y_{0}=4.0$ and coupled compound Poisson process and geometric Brownian motion noises, on the time interval $I=[0.0,5.0]$, based on $M=100$ sample paths for each fixed time step, with the target solution calculated with $262144$ points. The order of strong convergence is estimated to be $p=1.039$, with the 95% confidence interval $[1.0103,1.0681]$.
Figure 15. Order of convergence of the strong error of the Euler method for the toggle-switch model Equation 8.24, based on Table 6.
Figure 16. Evolution of a sample solution of the toggle-switch model Equation 8.24.
Figure 17. Noise sample paths for the toggle-switch model Equation 8.24.
8.3.6. Mechanical structure under random Earthquake-like seismic disturbances
Now we consider a simple mechanical structure model driven by a random disturbance in the form of a transport process, simulating seismic ground-motion excitations, inspired by the model in [9] (see also [30, Chapter 18] and [20] with this and other models).
@@ 8.3.6. Mechanical structure under random Earthquake-like seismic disturbances
$\ddot{X}_{t}+2\zeta_{0}\omega_{0}\dot{X}_{t}+\omega_{0}^{2}X_{t}=-\ddot{M}_{t}.$ (8.25)
@@
$M_{t}=\sum_{i=1}^{k}\gamma_{i}(t-\tau_{i})_{+}^{2}e^{-\delta_{i}(t-\tau_{i})}\cos(\omega_{i}(t-\tau_{i})),$ (8.26)
@@
Table 7. Mesh points (N), time steps (dt), strong error (error), and standard error (std err) of the Euler method for mechanical structure model under ground-shaking random excitations for each mesh resolution $N$, with initial condition $X_{0}=\mathbf{0}$ and transport process noise, on the time interval $I=[0.0,2.0]$, based on $M=100$ sample paths for each fixed time step, with the target solution calculated with $262144$ points. The order of strong convergence is estimated to be $p=1.147$, with the 95% confidence interval $[1.0595,1.2341]$.
Figure 18. Order of convergence of the strong error of the Euler method for the mechanical structure model Equation 8.25, based on Table 7.
Figure 19. Sample ground motion, ground velocity, and ground acceleration for the mechanical structure model Equation 8.25.
8.3.7. An actuarial risk model
@@
$\frac{\mathrm{d}X_{t}}{\mathrm{d}t}=R_{t}X_{t}+R_{t}(C_{t}+O_{t})+\nu O_{t}+\gamma.$ (8.27)
@@
As for the interest rate process $\{R_{t}\}_{t}$, there is a vast literature with models for it, see e.g. [10, Chapter 3] in particular Table 3.1. Here, we consider the Dothan model ([10, Section 3.2.2] of the aforementioned reference), which consists simply of a geometric Brownian motion process
@@
so that the equation Equation 8.27 for $\{X_{t}\}_{t}$ is a genuine random ODE.
@@
Table 8. Mesh points (N), time steps (dt), strong error (error), and standard error (std err) of the Euler method for a risk model for each mesh resolution $N$, with initial condition $X_{0}=1.0$ and coupled Ornstein-Uhlenbeck, geometric Brownian motion, and compound Poisson processes, on the time interval $I=[0.0,3.0]$, based on $M=400$ sample paths for each fixed time step, with the target solution calculated with $262144$ points. The order of strong convergence is estimated to be $p=0.976$, with the 95% confidence interval $[0.8938,1.0584]$.
Figure 20. Order of convergence of the strong error of the Euler method for the actuarial risk model Equation 8.27, based on Table 8.
Figure 21. Sample noises for the risk model Equation 8.27.
Figure 22. Sample surplus solution for the risk model Equation 8.27.
8.3.8. A random Fisher-KPP nonlinear PDE driven by boundary noise

### v8

@@ A.2.2. Fractional Brownian motion noise
$\begin{cases}\displaystyle\frac{\mathrm{d}X_{t}}{\mathrm{d}t}=-X_{t}+B^{H}_{t},\qquad 0\leq t\leq T,\\ \left.X_{t}\right|_{t=0}=X_{0},\end{cases}$ (A.9)
and where the noise $\{B^{H}_{t}\}_{t}$ is assumed to be a fractional Brownian motion (fBm) with Hurst parameter $0<H<1$. It turns out that, for $0<H<1/2$, the order of convergence is $H+1/2,$ while for $1/2\leq H<1,$ the order of convergence is 1. In any case, this is higher than the order $H$ estimated in previous works. The same seems to hold for a nonlinear dependency on the fBm, but the proof is more involved, depending on a fractional Itô formula (see [8, Theorem 4.2.6], [5, Theorem 4.1]and [38, Theorem 2.7.4]), based on the Wick Itô Skorohod (WIS) integral (see [8, Chapter 4]). A corresponding WIS isometry is also needed (see e.g. [8, Theorem 4.5.6]), involving Malliavin calculus and fractional derivatives. For these reasons, we leave the nonlinear case to a subsequent work and focus on this simple linear example, which suffices to illustrate the peculiarity of the dependence on $H$ of the order of convergence. For this linear equation, the proof of convergence is done rigorously below, with the framework developed in the main article.
We need to estimate the last term in Equation 6.3 of Proposition 6.1, involving the steps of the term f(t,x,y)=x+yf(t,x,y)=-x+y, which in this case reduce to
$f(s,X_{\tau^{N}(s)}^{N},Y_{s})-f(\tau^{N}(s),X_{\tau^{N}(s)}^{N},Y_{\tau^{N}(s)})=B^{H}_{s}-B^{H}_{\tau^{N}(s)},$ (A.10)
for $0\leq s\leq T$. There are several ways to represent an fBm (see e.g. [8, 38]). We use the formula (see [34, eq. (2.1)] or [8, eq. (1.1)])
$B^{H}_{t}=\frac{1}{\Gamma(H+1/2)}\left(\int_{-\infty}^{0}\left((t-s)^{H-1/2}-(-s)^{H-1/2}\right)\;\mathrm{d}W_{s}\right.\\ \left.+\int_{0}^{t}(t-s)^{H-1/2}\;\mathrm{d}W_{s}\right),$ (A.11)
where $\Gamma(\cdot)$ is the well-known Gamma function. For the step, Equation A.11 means that
$B^{H}_{s}-B^{H}_{\tau^{N}(s)}=\frac{1}{\Gamma(H+1/2)}\left(\int_{-\infty}^{\tau^{N}(s)}\left((s-\xi)^{H-1/2}-(\tau^{N}(s)-\xi)^{H-1/2}\right)\;\mathrm{d}W_{\xi}\right.\\ \left.+\int_{\tau^{N}(s)}^{s}(s-\xi)^{H-1/2}\;\mathrm{d}W_{\xi}\right).$ (A.12)
Then, using the stochastic Fubini Theorem to exchange the order of integration (see, again, [41, Section IV.6]),
$\int_{0}^{t_{j}}$ $\left(f(s,X_{\tau^{N}(s)}^{N},Y_{s})-f(\tau^{N}(s),X_{\tau^{N}(s)}^{N},Y_{\tau^{N}(s)})\right)\;\mathrm{d}s$ (A.13)
@@
$\frac{1}{\Gamma(H+1/2)}\Delta t_{N}\mathbb{E}\left[\left|\int_{-\infty}^{0}\int_{0}^{t_{j}}\left((s-\xi)^{H-1/2}-(\tau^{N}(s)-\xi)^{H-1/2}\right)\;\mathrm{d}s\;\mathrm{d}W_{\xi}\right|\right]\\ \leq c_{H}^{(1)}\Delta t_{N},$ (A.14)
@@
$\frac{1}{\Gamma(H+1/2)}\mathbb{E}\left[\left|\int_{0}^{t_{j}}\int_{\tau^{N}(\xi)+\Delta t_{N}}^{t_{j}}\left((s-\xi)^{H-1/2}-(\tau^{N}(s)-\xi)^{H-1/2}\right)\;\mathrm{d}s\;\mathrm{d}W_{\xi}\right|\right]\\ \leq c_{H}^{(2)}\Delta t_{N},$ (A.15)
@@
$\frac{1}{\Gamma(H+1/2)}\mathbb{E}\left[\left|\int_{0}^{t_{j}}\int_{\xi}^{\tau^{N}(\xi)+\Delta t_{N}}(s-\xi)^{H-1/2}\;\mathrm{d}s\;\mathrm{d}W_{\xi}\right|\right]\leq c_{H}^{(3)}\Delta t_{N}^{H+1/2},$ (A.16)
@@
Putting the three estimates Equation A.14, Equation A.15, Equation A.16 in Equation A.13 we find that
$\mathbb{E}\left[\left|\int_{0}^{t_{j}}\left(f(s,X_{\tau^{N}(s)}^{N},Y_{s})-f(\tau^{N}(s),X_{\tau^{N}(s)}^{N},Y_{\tau^{N}(s)})\right)\;\mathrm{d}s\right|\right]\\ \leq c_{H}^{(4)}\Delta t_{N}+c_{H}^{(3)}\Delta t_{N}^{H+1/2},$ (A.17)
where $c_{H}^{(4)}=c_{H}^{(1)}+c_{H}^{(2)}$. Using this estimate in Proposition 6.1 shows that the Euler method is of order $H+1/2$, when $0<H<1/2$, and is of order 1, when $1/2\leq H<1$, having in mind that $H=1/2$ corresponds to the classical Wiener process.
@@
Theorem A.1.
Consider the equation Equation A.9 where $\{B^{H}_{t}\}_{t}$ is a fractional Brownian motion (fBm) with Hurst parameter $0<H<1$. Suppose the initial condition $X_{0}$ satisfies
$\mathbb{E}[\|X_{0}\|^{2}]<\infty.$ (A.18)
Then, the Euler scheme for this initial value problem is of strong order $H+1/2$, for $0<H<1/2$, and is of order $1$, for $1/2\leq H<1$. More precisely,
$\max_{j=0,\ldots,N}\mathbb{E}\left[\left|X_{t_{j}}-X_{t_{j}}^{N}\right|\right]\leq c_{1}\Delta t_{N}+c_{2}\Delta t_{N}^{H+1/2},\qquad\forall N\in\mathbb{N},$ (A.19)
@@
The fractional Browninan motion noise term is generated with the $\mathcal{O}(N\log N)$ fast Fourier transform (FFT) method of Davies and Harte, as presented in [15] (see also [20, Section 14.4]). Table 4 shows the obtained convergence estimates, for a number of Hurst parameters, which is illustrated in Figure 12, matching the theoretical estimate of $p=\min\{H+1/2,1\}.$
@@
A.2.3. Population dynamics with harvest
Here, we consider a population dynamics problem modeled by a logistic equation with random coefficients, loosely inspired by [20, Section 15.2], with an extra term representing harvest:
$\frac{\mathrm{d}X_{t}}{\mathrm{d}t}=\Lambda_{t}X_{t}(1-\frac{X_{t}}{r})-\alpha H_{t}\frac{X_{t}}{r+X_{t}},$ (A.21)
@@ A.2.3. Population dynamics with harvest
$0<\gamma-\varepsilon\leq\Lambda_{t}\leq\gamma+\varepsilon<2\gamma,\qquad\forall t\geq 0.$
@@
The equation Equation A.21 becomes
@@
Table 5. Mesh points (N), time steps (dt), strong error (error), and standard error (std err) of the Euler method for population dynamics for each mesh resolution NN, with initial condition $X_{0}\sim\mathrm{Beta}(7.0,5.0)$ and gBm and step process noises, on the time interval $I=[0.0,1.0]$, based on $M=100$ sample paths for each fixed time step, with the target solution calculated with $262144$ points. The order of strong convergence is estimated to be $p=1.009$, with the 95% confidence interval $[0.9603,1.0571]$.
Figure 13. Order of convergence of the strong error of the Euler method for the population dynamics equation Equation A.21, based on Table 5.
Figure 14. An approximation sequence of a sample path solution of the population dynamics equation Equation A.21.
A.2.4. A toggle-switch model for gene expression
@@
$\begin{cases}\frac{\displaystyle\mathrm{d}X_{t}}{\displaystyle\mathrm{d}t}=\left(A_{t}+\frac{\displaystyle X_{t}^{4}}{\displaystyle a^{4}+X_{t}^{4}}\right)\left(\frac{\displaystyle b^{4}}{\displaystyle b^{4}+Y_{t}^{4}}\right)-\mu X_{t},\\ \frac{\displaystyle\mathrm{d}Y_{t}}{\displaystyle\mathrm{d}t}=\left(B_{t}+\frac{\displaystyle Y_{t}^{4}}{\displaystyle c^{4}+Y_{t}^{4}}\right)\left(\frac{\displaystyle d^{4}}{\displaystyle d^{4}+X_{t}^{4}}\right)-\nu Y_{t},\end{cases}$ (A.22)
@@
Table 6. Mesh points (N), time steps (dt), strong error (error), and standard error (std err) of the Euler method for a toggle-switch model of gene regulation for each mesh resolution NN, with initial condition $X_{0}=4.0,Y_{0}=4.0$ and coupled compound Poisson process and geometric Brownian motion noises, on the time interval $I=[0.0,5.0]$, based on $M=100$ sample paths for each fixed time step, with the target solution calculated with $262144$ points. The order of strong convergence is estimated to be $p=1.039$, with the 95% confidence interval $[1.0103,1.0681]$.
Figure 15. Order of convergence of the strong error of the Euler method for the toggle-switch model Equation A.22, based on Table 6.
Figure 16. Evolution of a sample solution of the toggle-switch model Equation A.22.
Figure 17. Noise sample paths for the toggle-switch model Equation A.22.
A.2.5. Mechanical structure under random Earthquake-like seismic disturbances
Now we consider a simple mechanical structure model driven by a random disturbance in the form of a transport process, simulating seismic ground-motion excitations, inspired by the model in [10] (see also [39, Chapter 18] and [24] with this and other models).
@@ A.2.5. Mechanical structure under random Earthquake-like seismic disturbances
$\ddot{X}_{t}+2\zeta_{0}\omega_{0}\dot{X}_{t}+\omega_{0}^{2}X_{t}=-\ddot{M}_{t}.$ (A.23)
@@
$M_{t}=\sum_{i=1}^{k}\gamma_{i}(t-\tau_{i})_{+}^{2}e^{-\delta_{i}(t-\tau_{i})}\cos(\omega_{i}(t-\tau_{i})),$ (A.24)
@@
Table 7. Mesh points (N), time steps (dt), strong error (error), and standard error (std err) of the Euler method for mechanical structure model under ground-shaking random excitations for each mesh resolution NN, with initial condition $X_{0}=\mathbf{0}$ and transport process noise, on the time interval $I=[0.0,2.0]$, based on $M=100$ sample paths for each fixed time step, with the target solution calculated with $262144$ points. The order of strong convergence is estimated to be $p=1.147$, with the 95% confidence interval $[1.0595,1.2341]$.
Figure 18. Order of convergence of the strong error of the Euler method for the mechanical structure model Equation A.23, based on Table 7.
Figure 19. Sample ground motion, ground velocity, and ground acceleration for the mechanical structure model Equation A.23.
A.2.6. An actuarial risk model
@@
$\frac{\mathrm{d}X_{t}}{\mathrm{d}t}=R_{t}X_{t}+R_{t}(C_{t}+O_{t})+\nu O_{t}+\gamma.$ (A.25)
@@
As for the interest rate process $\{R_{t}\}_{t}$, there is a vast literature with models for it, see e.g. [11, Chapter 3] in particular Table 3.1. Here, we consider the Dothan model ([11, Section 3.2.2] of the aforementioned reference), which consists simply of a geometric Brownian motion process
@@
so that the equation Equation A.25 for $\{X_{t}\}_{t}$ is a genuine random ODE.
@@
Table 8. Mesh points (N), time steps (dt), strong error (error), and standard error (std err) of the Euler method for a risk model for each mesh resolution NN, with initial condition $X_{0}=1.0$ and coupled Ornstein-Uhlenbeck, geometric Brownian motion, and compound Poisson processes, on the time interval $I=[0.0,3.0]$, based on $M=400$ sample paths for each fixed time step, with the target solution calculated with $262144$ points. The order of strong convergence is estimated to be $p=0.976$, with the 95% confidence interval $[0.8938,1.0584]$.
Figure 20. Order of convergence of the strong error of the Euler method for the actuarial risk model Equation A.25, based on Table 8.
Figure 21. Sample noises for the risk model Equation A.25.
Figure 22. Sample surplus solution for the risk model Equation A.25.
A.2.7. A random Fisher-KPP nonlinear PDE driven by boundary noise